Describe a first step for solving each system using elimination. Then solve each system.
3x+4y=-10
5x-2y=18

Answer by checkley71(8403)

3X+4Y=-10 & 5X-2Y=18 MULTIPLYING THE SECOND EQUATION BY 2 WE GET 10X-4Y=36 NOW ADD TO THE FIRST EQUATION WE GET
3X+4Y=-10
10X-4Y=36
-------------
13X=26 OR X=26/13 OR X=2 NOW SUBSTITUTING 2 FOR X IN EQUATION 1 WE GET
3*2+4Y=-10 OR 6+4Y=-10 OR 4Y=-16 OR Y=-16/4 OR Y=-4